OpenStudy (anonymous):

write an equation of the line that passes through (2,-7) and is parallel to the line x=5. Please show all steps!

OpenStudy (anonymous):

x=2 line

OpenStudy (anonymous):

Please show your work

OpenStudy (anonymous):

there is nothing in this very easy ...which is the line parallel to x=5 and passing through the point x=2 ...the only possibility is x=2 line

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

understood?

OpenStudy (anonymous):

kind of, i still need to show some sort of work of how i get the answer though

OpenStudy (anonymous):

ok wait a sec?

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

It would help if you used equations like y=mx+b in order to show me :)

OpenStudy (helder_edwin):

a line with equation x=5 has slope equal to \(\infty\) (becuase it is a vertical line) ok?

OpenStudy (anonymous):

look at the graph once vertical axis is x axis

OpenStudy (anonymous):

@helder_edwin ok...and?

OpenStudy (helder_edwin):

well this is not exactly correct but u can think of the slope as \[\large m=\frac{k}{0} \] where k is any constant.

OpenStudy (helder_edwin):

now the formula for the slope is \[\large m=\frac{y_1-y_2}{x_1-x_2} \] replacing m above and the point given u would get \[\large \frac{1}{0}=\frac{y-(-7)}{x-2}=\frac{y+7}{x-2} \] and from this \[\large 1(x-2)=0(y+7)\] or \[\large x-2=0 \qquad\text{or}\qquad x=2 \]

OpenStudy (anonymous):

in that sense literally by applying the line equation of a line with slope m and passes through a point (x*,y*) is y-y* = m(x-x*) following this (y-(-7)) =infinity ( x-2) (y+7)/infinity = x-2 but any number/infinity = 0 therefore x-2 =0 which implies x=2 is the line.

OpenStudy (helder_edwin):

got it @ineedhelp77 ??

OpenStudy (anonymous):

I'm still thinking...

OpenStudy (anonymous):

whats the problem?

OpenStudy (anonymous):

It's just that the way you guys solved the problem is waaaaaaaaaay different than the way I'm learning how to and my head is just spinning.

OpenStudy (helder_edwin):

take your time.

OpenStudy (anonymous):

if i were to solve this on paper, instead of writing infinity, would I write 0?

OpenStudy (anonymous):

lets solve it your way...what is the first step which you think?

OpenStudy (anonymous):

well, my teachers always told me to write y=mx+b first

OpenStudy (anonymous):

actually the equation of a line in most general form is not that .that is the problem

OpenStudy (anonymous):

No, but then I would usually take the same slope and then plug everything in

OpenStudy (anonymous):

y = mx + c is a simplified form of a line with some conditions

OpenStudy (anonymous):

i'll teach you the most general method if you wish which helps you solve any equation of a line

OpenStudy (anonymous):

That would be nice! I've never even seen y=mx+c before

OpenStudy (anonymous):

that is the same equation you wrote we always use c instead of b

OpenStudy (anonymous):

ok lets start

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

here, we always use b

OpenStudy (anonymous):

see first thing is to find slope ...do you know how to do this..if yes we'll go to next step

OpenStudy (anonymous):

yes i do know

OpenStudy (anonymous):

to remind you if m is the slope and we have two points (x1,y1) and (x2,y2) then m=(y2-y1)/(x2-x1) ...are you ok with this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then you can write (y2-y1) = m (x2-x1) ..ok?

OpenStudy (anonymous):

so is this how i start the problem? by writing this?

OpenStudy (anonymous):

wait a sec let me complete...you understood the previous step?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then for any general point (x,y) we can write the equation of a line by replacing (x2,y2) with (x,y)....then the equation becomes {(y-y1) = m (x-x1)} this is the equation of a line in most general form

OpenStudy (anonymous):

ok. so how do i use it for this problem?

OpenStudy (anonymous):

dont i need to find the slope first? how would i do that with only one point given?

OpenStudy (anonymous):

now first find slope m=(y2-y1)/(x2-x1) but as you know x=5 line value of x is constant then slope becomes infinity...understood?

OpenStudy (anonymous):

what should i write instead of infinity?

OpenStudy (anonymous):

you write infinity symbol ∞

OpenStudy (anonymous):

ok with it?

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

how would the distributing part work? its infinity(x-2)

OpenStudy (anonymous):

in given question the line which we need is parallel to x=5 then its slope should also be infinity ok?

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

now as you can see we have slope m and a point (2,-7)

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

by using the equation (y-(-7)) = infinity (x-2)

OpenStudy (anonymous):

okay i got that, but how would i distribute that

OpenStudy (anonymous):

you know (any number/infinity )= 0 ??

OpenStudy (anonymous):

is this understood?

OpenStudy (anonymous):

yeah, but arnt i supposed to multiply?

OpenStudy (anonymous):

you can take the infinity to the L.H.S and write (y+7)/infinity = (x-2)

OpenStudy (anonymous):

ok with this?

OpenStudy (anonymous):

LHS?

OpenStudy (anonymous):

left hand side :)

OpenStudy (anonymous):

oooh okayyy

OpenStudy (anonymous):

then what do you get??

OpenStudy (anonymous):

x=2!!! OMG THANKS SO MUCH

OpenStudy (anonymous):

You're the best thanks so much! :D

OpenStudy (anonymous):

if you wish i'll say something which you may follow

OpenStudy (anonymous):

sorry, i have to go now. its dinner time and i have to finish the rest of my homework. thanks anyway though!

OpenStudy (anonymous):

goodbye! and thanks again!

OpenStudy (anonymous):

ok but what i wanted to say was..don't blindly follow the procedure..use ur common sense by drawing graphs...just by thinking you could solve the above problem